# Federated Unlearning: A Survey on Methods, Design Guidelines, and Evaluation Metrics

### v1

@@ Abstract
Federated Learning (FL) enables collaborative training of a Machine Learning (ML) model across multiple parties, facilitating the preservation of users’ and institutions’ privacy by keeping data stored locally. Instead of centralizing raw data, FL exchanges locally refined model parameters to build a global model incrementally. While FL is more compliant with emerging regulations such as the European General Data Protection Regulation (GDPR), ensuring the right to be forgotten in this context — allowing FL participants to remove their data contributions from the learned model — remains unclear. In addition, it is recognized that malicious clients may inject backdoors into the global model through updates, e.g. to generate mispredictions on specially crafted data examples.
@@ I Introduction
To achieve high-performance levels suitable for real-world applications, Machine Learning (ML) and, in particular, Deep Learning (DL) models need to access large amounts of data. In numerous scenarios, users willingly furnish their data to service providers in return for enhanced service quality and improved user experience. Practical examples include the training of natural language processing models on datasets derived from Amazon reviews [1] or the development of micro-video recommendation systems leveraging user data from platforms such as TikTok and Instagram [2]. Data analytics hold the promise of delivering unprecedented value to companies and organizations, e.g., as pointed out by its key role in influencing election campaigns in the so-called Facebook-Cambridge Analytica scandal [3].
@@
Furthermore, the emergence of increasingly privacy-conscious data owners and the adoption of data regulations have influenced the perception of traditional centralized ML/DL, where the training corpus frequently contains user data, e.g., habits or preferences. In this context, Federated Learning (FL) is considered a promising solution, laying the foundation for privacy-preserving ML. In FL, data remains distributed among various sources or data owners, and a global model is iteratively built by aggregating the contributions of participants (often referred to as clients), e.g., model updates or gradients. Consequently, clients exchange ephemeral information, locally computed on their private data and only valid for the current state of the global ML model.
Nevertheless, even under the FL paradigm, how to guarantee the right to be forgotten is not clear. Indeed, the collaboratively learned model inevitably embeds the knowledge extracted from individual user data, potentially posing a risk of leaking sensitive information [8] and vulnerability to membership attacks [9], i.e., the ability to infer whether a certain data sample was part of the training set. These considerations raise the following question: how should the FL system react when a client exercises its right to be forgotten at runtime?
@@ I-A Related Surveys
The interest in FU has been rapidly increasing. This trend is supported by the numerous research papers that have recently introduced novel FU algorithms. Most published works primarily focus on MU [6, 10, 13, 14], while FU is only covered from a general perspective with a limited amount of references.
By delving into distinct aspects of FU, a few other surveys have been proposed very recently. Wang et al. [11] review some studies on FU by mainly concentrating on the security and privacy dimensions of FU and by paying special attention to membership inference attacks and corresponding countermeasures. Liu et al. [12] present a comprehensive survey that highlights challenges, methods, and promising research directions in the field. The authors classify FU algorithms based on who initiates the unlearning and what needs to be forgotten. Additionally, their survey discusses optimization strategies for general FL. Similarly, Yang et al. [10] direct their attention toward challenges and future research directions, but by adopting a high-level perspective and without providing in-depth technical details. The authors categorize works according to the objectives of FU, with limited practical and implementation insights about the surveyed solutions. To provide a clear and concise comparison, Table I compares this paper with the other few surveys already conducted in the general field of MU.
@@ I-B Contributions
Specifically, to be highly accessible, we first present the fundamental concepts and principles of FU, encompassing objectives, motivations, and challenges. Then, we describe the primary emerging guidelines for the design of efficient FU algorithms, by including a comprehensive analysis of the metrics to assess unlearning - a critical concern within federated scenarios. Notably, we complement our reported analysis with a series of experiments that offer tangible insights into the necessity of FU. Moreover, we conduct an extensive review of the existing FU literature, categorizing papers through a novel taxonomy based on their objectives and the metrics employed. Finally, we identify open problems and discuss potential future research directions. We believe that this survey serves as a valuable resource for researchers seeking a clear understanding of FU and the recent related research advancements and as a practical guide for efficiently designing and implementing novel FU solutions. The paper contributions can be summarized as follows:
@@ I-C Organization
The remainder of this paper is organized as follows. Section II introduces the needed background about MU and FL. In Section III, we formalize the objectives of FU, we provide qualitative experimental evidence that motivates the research in the field, and we clarify the technical motivations why FU is more challenging than regular MU. In Section IV, we offer a detailed analysis of the requirements that FU algorithms have to meet and the associated assessment metrics. Furthermore, Section V presents an in-depth exploration of all the surveyed solutions, by highlighting the rationale behind their algorithm design and implementation choices. Section VI and VII respectively summarize the key lessons learned and the most promising directions for future work in FU, while Section VIII concludes the manuscript.
@@ II-A Machine Unlearning
Considering a set of model weights $w$ trained on a dataset $D$, the purpose of MU is to remove the influence of a specific $D$’s subset on $w$.
Let $D_{u}$ be the subset of $D$ to be forgotten, and let $D_{r}$ be its complement, i.e. $D_{u}$:$=D_{r}^{\complement}$ $=$ $D/D_{r}$. We will refer to $D_{u}$ as the forgetting dataset and to $D_{r}$ as the retain dataset. Let $w_{r}$ be a set of model weights trained only on $D_{r}$, usually referred to as the retrained model. The goal of an MU algorithm is to efficiently obtain $w_{u}$, a sanitized version of $w$, which is indistinguishable or approximately indistinguishable from $w_{r}$ [15, 6]. Two sets of model weights are indistinguishable if an attacker with access to $w_{u}$ cannot design a function $g(\cdot)$ to extract information specific to $D_{u}$. Considering, for example, a Kullback-Leibler (KL) divergence as a measure for similarity among distributions (distributions of weights in this case), the unlearned model weights $w_{u}$ and the retrained model weights $w_{r}$ are formally indistinguishable if their KL divergence is zero [15]:
$KL(g(w_{u}),g(w_{r}))=0$ (1)
with $w_{u}=\mathcal{U}(w)$ and $\mathcal{U}(\cdot)$ representing an unlearning algorithm. Similarly, the unlearned model weights $w_{u}$ and the retrained model weights $w_{r}$ are approximately (or $\epsilon$-indistinguishable) if their KL divergence is lower than a small threshold value $\epsilon$,
$KL(g(w_{u}),g(w_{r}))<\epsilon.$ (2)
@@
It is worth noting that MU mechanisms are not only needed to meet regulations about individuals’ privacy, such as GDPR and CCPA, but also represent a security tool for the provider of the ML service. Training data may include natural or malicious outliers, which can be detrimental to the model performances or can introduce potential security threats (e.g., model backdoor). Under this perspective, MU algorithms represent a mechanism to selectively remove the influence of such a subset of threatening data without discarding the useful knowledge already acquired.
@@ II-B Federated Learning
Federated Averaging [16], also known as FedAvg, serves as the foundational algorithm for FL and can be described as a synchronous protocol that adopts a client-server paradigm [17]. FedAvg proceeds in rounds and heuristically tries to solve the optimization problem in Eq. 3. During each round, the server distributes the current global model to a random selection of available clients among the entire pool of learners, which fine-tune the received model parameters on their local data. Then, the activated clients communicate a model update to the central server (i.e., the difference between the locally fine-tuned and the received model parameters). Once the server has gathered such model updates, a weighted-average aggregation is performed, and the resultant averaged update is applied to the current global model. The initial formulation of FedAvg applies the resultant aggregated update by just summing it to the server model. However, aggregated updates can be seen as pseudo-gradients, and can be applied by using an optimizer of choice, i.e., different from regular Stochastic Gradient Descent (SGD). Under this perspective, vanilla FedAvg employs SGD with a learning rate set to 1 as server-side optimizer [18].
@@ II-C Catastrophic Forgetting and Differential Privacy
In this subsection, we introduce two concepts, i.e. catastrophic forgetting and Differential Privacy, that are similar to unlearning algorithms in their effects while being substantially different in principle.
@@
Catastrophic forgetting. In the field of Continual Learning (CL) [21], a key challenge is catastrophic forgetting. This occurs when a neural network, while adapting to new tasks, tends to disrupt the representations learned to solve the previous downstream tasks [22]. Recent studies have demonstrated that catastrophic forgetting also happens in FL training due to the round-by-round shift of data distributions. The pools of selected participants can change at each round, potentially including clients with skewed data distribution with respect to the global data distribution. This can lead the global model to specialize mainly on the specific per-round data distribution of activated clients while struggling to generalize overall [23, 24, 25].
@@
Catastrophic forgetting may seem a natural method for unlearning, i.e., the global model forgets the specific contributions of a client’s data as soon as that client is not included in FL rounds anymore. However, it has been demonstrated [26] (see also Subsection III-B) that even though a participant leaves the training, the global model still retains contributions learned from its data. Consequently, relying on catastrophic forgetting cannot be considered a reliable strategy. Moreover, when unlearning methods are applied to an ML model, there is the risk that catastrophic forgetting is induced by the unlearning, i.e. the unlearning algorithm also removes to-retain knowledge as a side effect.
@@
Differential Privacy. Within the context of ML, it is of paramount importance that the public release of model parameters does not leak too much sensitive information about the original training data, even when the adversary can have access to auxiliary information. Differential Privacy (DP) [27, 28] tries to solve this apparently paradoxical requirement of extracting useful information from a population while guaranteeing to learn nothing about an individual. To this end, DP promises that the inclusion, or exclusion, of an individual’s data in a private database does not essentially affect the outcome of any analysis conducted on that database. Therefore, an individual should be almost indifferent between disclosing her data and not [^1] If the inclusion (or exclusion) of an individual’s data did not affect the outcome at all, it would mean that the individual’s data is useless.. In a nutshell, DP is a mathematically rigorous formalization of this promise, capable of quantifying the qualitative indications of too much, essentially, and almost thanks to a privacy measure, a parameter $\delta$, referred to as privacy budget. When applied to ML/DL models, DP injects $\epsilon$-calibrated noise, e.g. drawn from a Gaussian distribution, to perturb model updates or gradients, at the cost of possible performance degradation (e.g., [29, 30, 31]), with $\epsilon$ being the parameter that tunes the noise magnitude, the smaller the $\epsilon$ the stronger the noise. In FL processes, DP should be employed from the beginning, hence it cannot be seamlessly integrated into ongoing training; furthermore, DP is not helpful when there is the need to remove harmful contributions from malicious clients. From these observations, FU can be considered an orthogonal tool with respect to DP.
@@
Figure 2: Federated unlearning overview. When the client U requests to be forgotten, the FL server erases its contributions from the global model $w_{t}$, generating a new version $w_{t}^{\bar{u}}$.
@@ III Federated Unlearning
Suppose that after $t$ rounds of FL, a given client $u$ submits an unlearning request for a subset $S_{u}$ of its private data $D_{u}$ and an unlearning algorithm $\mathcal{U}$ is executed to fulfill the request. An unlearning algorithm $\mathcal{U}$ is applied to the global model $w_{t}$, which has been trained including client $u$ in the pool of clients, generating an unlearned global model $w^{\bar{u}}_{t}$, i.e. $w^{\bar{u}}_{t}=\mathcal{U}(w_{t})$. $\mathcal{U}(\cdot)$ can be defined as a function that ensures that $w^{\bar{u}}_{t}$ exhibits performances indistinguishable or approximately indistinguishable from a global model $w^{r}_{t}$ trained without the contribution of $S_{u}$.
FU can have different objectives (sample unlearning, class unlearning, and client unlearning), as we detail in Subsection III-A. In Subsection III-B we qualitatively and experimentally motivate the need for unlearning even though FL does not require direct access to raw data. In fact, the weights of the global model being trained round by round may leak information about clients’ data. As we describe in Subsection III-C, the architectural design of FL introduces a more challenging setting than regular MU, due to the decentralization and inscrutability of data, and the iterative and stochastic nature of the FL process. In Subsection IV-A we outline the requirements that FU algorithms should meet, while in the last Subsection IV-B we outline and describe the metrics being used in the related literature to assess that an FU mechanism fulfills the previously introduced requirements.
@@ III-A Objectives
As depicted in Fig. 3, FU aims to achieve one or more of the following unlearning objectives [32]:
Sample Unlearning. Sample unlearning aims to remove the contribution of specific data samples from the trained model. In this case, the client that requests the unlearning would like to erase the contribution of a subset of its data, i.e. $S_{u}\subset D_{u}$.
Class Unlearning. Class unlearning aims to remove the contribution of all the data samples that belong to a certain class $c$ across clients, i.e. to remove the contribution of $C=\bigcup_{k\in K}S_{k}$ with $S_{k}=\{x_{i},c\}^{n^{c}_{k}}_{i}$, with $n^{c}_{k}$ the number of local samples at client $k$ labeled with class $c$. Ideally, the unlearned model, if provided with samples belonging to a removed class, should produce an outcome at random between the remaining classes.
Client Unlearning. Client unlearning is peculiar to federated settings, and it relates to the right to be forgotten of clients. In this case, the client that requests the unlearning would like to erase the contribution of its entire local dataset, i.e. $S_{u}=D_{u}$. Client Unlearning can be seen as applying Sample Unlearning on all the samples that a specific client held at the time of local training. Similarly, Client Unlearning can also be a special case of class unlearning when a single client $u$ is the only one holding all the samples for class $c$ across the federation, i.e., $C=S_{u}=\{x_{i},c\}^{n_{u}}_{i}$ and $S_{u}=D_{u}$.

### v3

@@ I Introduction
To achieve high-performance levels suitable for real-world applications, Machine Learning (ML) and, in particular, Deep Learning (DL) models need to access large amounts of data. In numerous scenarios, users willingly furnish their data to service providers for enhanced service quality and improved user experience. Practical examples include the training of natural language processing models on datasets derived from Amazon reviews [1] or the development of micro-video recommendation systems leveraging user data from platforms such as TikTok and Instagram [2]. Data analytics promise to deliver unprecedented value to companies and organizations, e.g., as pointed out by its key role in influencing election campaigns in the so-called Facebook-Cambridge Analytica scandal [3].
@@
Furthermore, the emergence of increasingly privacy-conscious data owners and the adoption of data regulations have influenced the perception of traditional centralized ML/DL, where the training corpus frequently contains user data, e.g., habits or preferences. In this context, Federated Learning (FL) is considered a promising solution, laying the foundation for privacy-preserving ML. In FL, data remains distributed among various sources or data owners, and a global model is iteratively built by aggregating the contributions of participants (often referred to as clients), e.g., model updates or gradients. Consequently, clients exchange ephemeral information locally computed on their private data, which is only valid for the current state of the global ML model.
Nevertheless, even under the FL paradigm, how to guarantee the right to be forgotten is unclear. Indeed, the collaboratively learned model inevitably embeds the knowledge extracted from individual user data, potentially posing a risk of leaking sensitive information [8] and vulnerability to membership attacks [9], i.e., the ability to infer whether a certain data sample was part of the training set. These considerations raise the following question: how should the FL system react when a client exercises their right to be forgotten at runtime?
@@ I-A Related Surveys
Most published surveys primarily focus on MU [6, 11, 14, 15, 16], while FU is only covered from a general perspective with a limited amount of references.
By delving into distinct aspects of FU, a few other surveys have been proposed very recently. Wang et al. [12] review some studies on FU by mainly concentrating on the security and privacy dimensions of FU and by paying special attention to membership inference attacks and corresponding countermeasures. Liu et al. [13] present a comprehensive survey that highlights challenges, methods, and promising research directions in the field. The authors classify FU algorithms based on who initiates the unlearning and what needs to be forgotten. Methods are classified as either passive or active, depending on the client’s role in the forgetting process. In addition, the paper highlights some open problems specific to FU, such as its emerging threats and the need to reach fairness and explainability. Similarly, Yang et al. [11] direct their attention toward challenges and future research directions, but by adopting a high-level perspective and without providing in-depth technical details. The authors categorize works according to the objectives of FU, with limited practical and implementation insights about the surveyed solutions. To provide a clear and concise comparison, Table I compares this paper with the other few surveys already conducted in the general field of FU.
@@ I-B Contributions
Specifically, to be highly accessible, we first present the fundamental concepts and principles of FU, encompassing objectives, motivations, and challenges. Then, we describe the primary emerging guidelines for the design of efficient FU algorithms, by including a comprehensive analysis of the metrics to assess unlearning - a critical concern within federated scenarios. Notably, we complement our reported analysis with a series of experiments that offer tangible insights into the necessity of FU. Moreover, we conduct an extensive review of the existing FU literature, categorizing papers through a novel taxonomy based on their objectives and the metrics employed. Finally, we identify open problems and discuss potential future research directions. We believe that this survey serves as a valuable resource for researchers seeking a clear understanding of FU and the recent related research advancements and as a practical guide for efficiently designing and implementing novel FU solutions. The paper’s contributions can be summarized as follows:
@@ I-C Organization
The remainder of this paper is organized as follows. Section II introduces the needed background about MU and FL. In Section III, we formalize the objectives of FU, we provide qualitative experimental evidence that motivates the research in the field, and we clarify the technical motivations why FU is more challenging than regular MU. In Section IV, we offer a detailed analysis of the requirements that FU algorithms have to meet and the associated assessment metrics. Furthermore, Section V presents an in-depth exploration of all the surveyed solutions, by highlighting the rationale behind their algorithm design and implementation choices. Section VI and VII respectively summarize the key lessons learned and the most promising directions for future work in FU. Finally, Section VIII concludes the manuscript. The structure of this survey is highlighted in Figure 1, while Table II reports the notation adopted throughout the paper.
@@ II-A Machine Unlearning
Considering a set of model weights $w$ trained on a dataset $D$, the purpose of MU is to remove the influence of a specific $D$’s subset on $w$. Let $D_{u}$ be the subset of $D$ to be forgotten, and let $D_{r}$ be its complement, i.e., $D_{u}$:$=D_{r}^{\complement}$ $=$ $D/D_{r}$. We will refer to $D_{u}$ as the forgetting dataset and to $D_{r}$ as the retain dataset. Let $w^{r}$ be a set of model weights trained only on $D_{r}$, usually referred to as the retrained model. The goal of an MU algorithm is to efficiently obtain $w^{\bar{u}}$, a sanitized version of $w$, which is indistinguishable or approximately indistinguishable from $w^{r}$ [17, 6]. Two sets of model weights are indistinguishable if an attacker with access to $w^{\bar{u}}$ cannot design a function $g(\cdot)$ to extract information specific to $D_{u}$. Considering, for example, a Kullback-Leibler (KL) divergence as a measure for similarity among distributions (distributions of weights in this case), the unlearned model weights $w^{\bar{u}}$ and the retrained model weights $w^{r}$ are formally indistinguishable if their KL divergence is zero [17]:
$KL(g(w^{\bar{u}}),g(w^{r}))=0$ (1)
with $w^{\bar{u}}=\mathcal{U}(w)$ and $\mathcal{U}(\cdot)$ representing an unlearning algorithm. Similarly, the unlearned model weights $w^{\bar{u}}$ and the retrained model weights $w^{r}$ are approximately (or $\epsilon$-indistinguishable) if their KL divergence is lower than a small threshold value $\epsilon$,
$KL(g(w^{\bar{u}}),g(w^{r}))<\epsilon.$ (2)
@@
It is worth noting that MU mechanisms are not only needed to meet regulations about individuals’ privacy, such as GDPR and CCPA, but also represent a security tool for the provider of the ML service. Training data may include natural or malicious outliers, which can harm the model performances or introduce potential security threats (e.g., model backdoor). Under this perspective, MU algorithms represent a mechanism to selectively remove the influence of such a subset of threatening data without discarding the useful knowledge already acquired.
@@ II-B Federated Learning
Federated Averaging [18], also known as FedAvg, serves as the foundational algorithm for FL and can be described as a synchronous protocol that adopts a client-server paradigm [19]. FedAvg proceeds in rounds and heuristically tries to solve the optimization problem in Eq. 3. During each round, the server distributes the current global model to a random selection of available clients, which fine-tune the received model parameters on local data. Then, the selected clients communicate a model update to the server (i.e., the difference between the locally fine-tuned and the received model parameters). Once the server has gathered such model updates, a weighted-average aggregation is performed, and the resultant averaged update is applied to the current global model. The initial formulation of FedAvg applies the resultant aggregated update by just summing it to the server model. However, aggregated updates can be seen as pseudo-gradients, and can be applied by using an optimizer of choice, i.e., different from regular Stochastic Gradient Descent (SGD) [20].
@@ II-C Catastrophic Forgetting and Differential Privacy
In this subsection, we introduce catastrophic forgetting and Differential Privacy (DP), two concepts correlated with FU. In particular, we clarify why these concepts are orthogonal to unlearning.
@@
Catastrophic forgetting. In the field of Continual Learning (CL) [23], a key challenge is catastrophic forgetting. This occurs when a neural network while adapting to new tasks, tends to disrupt the representations learned to solve the previous downstream tasks [24]. Recent studies have demonstrated that catastrophic forgetting also happens in FL training due to the round-by-round shift of data distributions. The pools of selected participants can change at each round, potentially including clients with skewed data distribution concerning the global data distribution. This can lead the global model to specialize mainly on the specific per-round data distribution of activated clients while struggling to generalize overall [25, 26, 27].
@@
Catastrophic forgetting may seem a natural method for unlearning, i.e., the global model forgets the specific contributions of a client’s data as soon as th at client is no longer included in FL rounds. However, it has been demonstrated [28] (see also Subsection III-B) that even though a participant leaves the training, the global model still retains contributions learned from its data. Consequently, relying on catastrophic forgetting cannot be considered a reliable strategy. Moreover, when unlearning methods are applied to an ML model, there is the risk that catastrophic forgetting is induced by the unlearning, i.e., the unlearning algorithm also removes to-retain knowledge as a side effect.
@@
Differential Privacy. Within the context of ML, it is crucial that the public release of model parameters does not leak too much sensitive information about the original training data, even if an adversary can have access to auxiliary information. DP [29, 30] addresses this challenge by ensuring that the inclusion, or exclusion, of an individual’s data in a private database does not essentially affect the outcome of any analysis conducted on that database. Therefore, an individual should be almost indifferent between disclosing her data and not [^1] If the inclusion (or exclusion) of an individual’s data did not affect the outcome at all, it would mean that the individual’s data is useless.. In a nutshell, DP is a rigorous mathematical framework to quantify privacy guarantees, formalizing the vague terms of too much, essentially, and almost thanks to a privacy measure called privacy budget, denoted as $\delta$. When applied to ML/DL models, DP injects $\epsilon$-calibrated noise, e.g., drawn from a Gaussian distribution, to perturb model updates or gradients. This noise introduces a trade-off, as it can degrade model performance (e.g., [31, 32, 33]), with $\epsilon$ being the parameter that tunes the noise magnitude, the smaller the $\epsilon$ the stronger the noise. In FL processes, DP should be employed from the beginning and it cannot be seamlessly integrated into ongoing training. DP can be considered an orthogonal tool for unlearning since it masks the participation of individuals, reducing the vulnerability to membership attacks. On the other hand, DP does not address the need to remove harmful contributions from malicious clients.
@@
Figure 2: FU overview. When the client $u$ requests to be forgotten, the FL server erases its contributions from the global model $w_{t}$, generating a new version $w_{t}^{\bar{u}}$.
@@ III Federated Unlearning
Suppose that after $t$ rounds of FL, a given client $u$ submits an unlearning request for a subset $S_{u}$ of its private data $D_{u}$ and an unlearning algorithm $\mathcal{U}$ is executed to fulfill the request. An unlearning algorithm $\mathcal{U}$ is applied to the global model $w_{t}$, which has been trained including client $u$ in the pool of clients, generating an unlearned global model $w^{\bar{u}}_{t}$, i.e., $w^{\bar{u}}_{t}=\mathcal{U}(w_{t})$. $\mathcal{U}(\cdot)$ can be defined as a function that ensures that $w^{\bar{u}}_{t}$ exhibits performances indistinguishable or approximately indistinguishable from a global model $w^{r}_{t}$ trained without the contribution of $S_{u}$. Figure 2 offers an illustrative overview of FU.
FU can have different objectives (sample unlearning, class unlearning, and client unlearning), as we detail in Subsection III-A. In Subsection III-B, we qualitatively and experimentally motivate the need for unlearning even though FL does not require direct access to raw data. In fact, the weights of the global model being trained round-by-round may leak information about clients’ data. As we describe in Subsection III-C, the architectural design of FL introduces a more challenging setting than regular MU, due to the decentralization and inscrutability of data, and the iterative and stochastic nature of the FL process. In Subsection IV-A, we present the requirements that FU algorithms should meet. Finally, Subsection IV-B outlines and describes the metrics being used in the related literature to assess whether an FU mechanism fulfills the previously introduced requirements.
@@ III-A Objectives
As depicted in Figure 3, FU aims to achieve one or more of the following unlearning objectives [34]:
Sample Unlearning. Sample unlearning aims to remove the contribution of specific data samples from the trained model. In this case, the client that requests the unlearning would like to erase the contribution of a subset of its data, i.e., $S_{u}\subset D_{u}$.
Class Unlearning. Class unlearning aims to remove the contribution of all the data samples that belong to a certain class $c$ across clients, i.e., to remove the contribution of $C=\bigcup_{k\in K}S_{k}$ with $S_{k}=\{x_{i},c\}^{n^{c}_{k}}_{i}$, with $n^{c}_{k}$ the number of local samples at client $k$ labeled with class $c$. Ideally, if provided with samples belonging to a removed class, the unlearned model should produce an outcome at random between the remaining classes.
Client Unlearning. Client unlearning is peculiar to federated settings, and it relates to the right to be forgotten of clients. In this case, the client that requests the unlearning would like to erase the contribution of its entire local dataset, i.e., $S_{u}=D_{u}$. Client Unlearning can be seen as applying Sample Unlearning on all the samples that a specific client held at the time of local training. Similarly, Client Unlearning can also be a special case of class unlearning when a single client $u$ is the only one holding all the samples for class $c$ across the federation, i.e., $C=S_{u}=\{x_{i},c\}^{n_{u}}_{i}$ and $S_{u}=D_{u}$.